OpenStudy (anonymous):

Find the next three terms of the sequence. Then write a rule for the sequence. 648, 216, 72, 24 I don't really understand this question

OpenStudy (ankit042):

it is a GP with a = 648 and r = 1/3

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

@ankit042 can you explain please?

OpenStudy (ankit042):

have you heard about Geometrical progression?

OpenStudy (ankit042):

it can be written as a ar ar^2 ar^3......ar^(n-1)

OpenStudy (anonymous):

@ankit042 i have "heard", i do not know

OpenStudy (ankit042):

sorry I can't teach you the basics to solve this question you have to find next three terms which will be 24/3=8 8/3 and so on

OpenStudy (anonymous):

@ankit042 Okay, thanks for your help. Would you happen to know how to right a rule for a sequence/what is a rule?

OpenStudy (ankit042):

I think it will be a generic term expressed in n so for your series lets call it S. S(n) = ar(n-1) where a =648,r=1/3 substitute these values and you get your rule

OpenStudy (anonymous):

@ankit042 I'm so sorry. This all complete gibberish to me. Math is not my strong point. I have no idea what you're talking about.